OpenStudy (anonymous):

How can I calculate the gravity if I use a toy car to run down a slope? Also if I change the distance of the starting point of the toy car on the same slope, what effect would it have?

OpenStudy (rane):

it depends on then wht measurements u r given

OpenStudy (anonymous):

Say the slope is 45 degrees from the ground. Slope is 1 meter. Height is 0.5 meters. (I made these measurements up). Are you sure the measurements matter? Isn't there a formula to find the gravity?

OpenStudy (anonymous):

I was told to use the formula \[S=UT+\frac{ 1 }{ 2 }AT ^{2}\] Where A is: \[A = g \cos \theta \] But the UT cancels out since the value of U = 0 So it would be: \[S=\frac{ 1 }{ 2 }g \cos \theta T ^{2}\] This is what I was told to use but i'm not sure how this helps me find the gravity.

OpenStudy (anonymous):

\[\frac{2 \Delta S}{\cos(\theta) \Delta t^{2}} = g\]

OpenStudy (anonymous):

What does the triangle mean? Sorry haven't been taught this haha. And if I change the distance of the starting point of the toy car would that effect anything?

OpenStudy (anonymous):

The triangle is a capital delta and it means a change in, so change in S or change in t

OpenStudy (anonymous):

Let's say the slope is 45 degrees. Height of the slope is 0.5m. The slope is 2m. I'll roll a ball down it and it takes 1.5 seconds to reach the bottom. To find the gravity would I use the formula Eleven17 mentioned and go. \[\frac{ 2x2 }{ \cos 45x1.5^{2} } = g\]

OpenStudy (rane):

u can

OpenStudy (anonymous):

What is that formula called? Does it have a particular name?

OpenStudy (rane):

in phys formula's dont have names bt they r known by wht we have to find using that formula

OpenStudy (anonymous):

Alright thanks guys really helped a lot. But one more thing. Would a good independent variable for my experiment of finding gravity be changing the value of S? I just have to calculate the gravity in my suburb.